OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x - 3y - 3z = -13 x + 2y - z = -1 2x + 3y + 2z = 0 How do I do this and what do you get for an answer?

OpenStudy (anonymous):

multiply equation 2 by 2: 2x + 4y - 2z = -2 add to no. 3 2x + 3y + 2z = 0 4x + 7y = -2......................(1) multiply second by 3 3x + 6y - 3z = -3 subtract first equation from this one x - 3y - 3z = -13 2x + 9y = 10.......................(2) multiply (2) by 2 4x + 18y = 20 subtract the above from (1): -11y = -22 y = 2 substitute y = 2 in equation (2): 2x + 18 = 10 x = -4 now substitute for x and y in first of original equatios x - 3y - 3z = -13 -4 - 6 - 3z = -13 3z = 3 z = 1 so x = -4,y=2, z = 1 check results in 3rd equation: 2(-4) + 3(2) + 2(1) = 0 ok

OpenStudy (anonymous):

phew!! - i'm a little tired!!

OpenStudy (anonymous):

is that what you wanted?

OpenStudy (anonymous):

Yep :] Thanks! Haha.